OpenStudy (summersnow8):

John stands on the edge of a deck that is 40.0 m above the ground and throws a rock straight up that reaches a height of 15.0 m above the deck. What is the initial speed of the rock in m/s? which i solved correctly and it is 17.1 Assuming the rock in the previous question misses the deck on the way down, how fast (in m/s) will the rock be moving when it hits the ground? @ybarrap @lightgrav

OpenStudy (summersnow8):

i used the equation \[2as = V _{f}^{2} - V{i}^{2}\]

OpenStudy (anonymous):

so, use that correct answer in the same statement (equation) as before, but solve it for v_f ... remember that (now) a is in the same direction as s (both downward) so 2as is positive.

OpenStudy (anonymous):

(or, you could start with v_i = 0, but watch the rock fall 55m

OpenStudy (summersnow8):

so, \[2 (-9.8)(-40) = Vf ^{2} - (17.1)^{2}\]

OpenStudy (anonymous):

looks right so far

OpenStudy (summersnow8):

so vf would be 32.8?

OpenStudy (anonymous):

yes, 55m is almost 4x the upward 15m (so v should be almost double)

OpenStudy (ybarrap):

You got it! Another way to look at this - $$ V_f^2 = V_i^2 + 2gs $$ Where \(V_i=0\), the speed of the rock at the highest point. \(s\) is is the distance the rock is above the deck plus the height the deck is above the ground: 15+40=55. Now you have everything to solve. You've got the right approach either way!

OpenStudy (summersnow8):

so is 32.8 the answer? i was told to use the equation i listed

OpenStudy (anonymous):

His eq'n is the same as yours, just re-arranged for v_f

OpenStudy (ybarrap):

Yep! $$ V_f^2 = V_i^2 + 2gs=0-2\times9.81\times55\\ \implies V_f=32.8~m/s $$

OpenStudy (ybarrap):

*Negative, because rock is going down: $$ V_f^2 = V_i^2 + 2gs=0-2\times9.81\times55\\ \implies V_f=\color{red}{-}32.8~m/s $$

OpenStudy (summersnow8):

..... negative?

OpenStudy (ybarrap):

Just indicates direction, downward. Up is positive

OpenStudy (anonymous):

the same direction as the acceleration.

OpenStudy (summersnow8):

it says it's wrong

OpenStudy (anonymous):

ooh, they asked for _speed_ (how fast) not velocity :-(

OpenStudy (ybarrap):

speed is magnitude, velocity includes directional information. So speed would just be 32.8 m/s